satellite73 (satellite73):

can someone help with the laurant expansion of \[\frac{1+z}{z}\] at 0? i get \[2+\sum_{k=1}^{\infty}(-1)^k (z-1)^k\] but i need the expansion for all z

OpenStudy (jamesj):

Do you mean the Laurant expansion at z = 1 ?

OpenStudy (anonymous):

maybe i am just being daft. maybe it is only \[\frac{1}{z}+1\] and i am thinking way too hard?

OpenStudy (jamesj):

At z = 0, the expansion is just 1/z + 1, yes.

OpenStudy (anonymous):

ok i get the dope slap award. i was trying to do it as if it was a geometric series, but now i am embarrassed.

OpenStudy (jamesj):

live and learn. Or calculate and learn.

OpenStudy (anonymous):

well here is one that i am actually going to have to think about. expand \[\frac{z}{z^2+1}\] about i on \[|z-i|<2\] any suggestions?

OpenStudy (anonymous):

partial fractions?

OpenStudy (jamesj):

\[ \frac{z}{z^2 + 1} = \frac{1}{z-i} \frac{z}{z+i} \] So if you have an expansion of the z/(z+i) term, you you just shift all the terms down by one power of (z-i). Now, for the z/(z+i) itself, you can start differentiating and evaluating at i, or write it as \[ \frac{z}{z+i} = 1 - \frac{i}{z+i} \] and start differentiating and evaluating that, which is very manageable.

OpenStudy (anonymous):

ok thank you so much.